OpenStudy (anonymous):

(5.8 x 10^6) (6.1 x 10^3) Multiply Write using scientific notation

OpenStudy (anonymous):

multiply 5.8 by 6.1, and then multiply this by 10^9

OpenStudy (callisto):

(5.8 x 10^6) (6.1 x 10^3) = (5.8 x 6.1) x10^(6+3) = 35.38 x10^9 = ??? Can you do the last step?

Directrix (directrix):

Following from above ...

OpenStudy (anonymous):

3.538 x 10^4

OpenStudy (anonymous):

If you look at the link directrix gave... you will see the answer... :D